OpenStudy (anonymous):

What is the vertex for the function? (You will want to find the axis of symmetry first.) Write as an ordered pair, leaving no spaces, such as (1,-4). y = -3x2 + 6x

OpenStudy (anonymous):

axis of symmetry is x=-b/(2a)=-6/(2*-3)=1 plug in 1 to y=-3x^2+6x to get your y-coordinate for the vertex

OpenStudy (anonymous):

Can you further explain?

OpenStudy (anonymous):

if you have a parabola in the form ax^2+bx+c the axis of symmetry is: x=-b/(2a)

OpenStudy (anonymous):

the vertex is at: \[(\frac{-b}{2a}, f(\frac{-b}{2a}))\]

OpenStudy (anonymous):

you can also rewrite this to a form you might more easily recognize: \[a(x-h)^2+k\] where (h,k) is the vertex y=-3x^2+6x y=-3(x^2-2x) y=-3(x^2-2x+1-1) y=-3(x-1)^2-3

OpenStudy (anonymous):

sorry, last line should be: y=-3(x-1)^2+3

OpenStudy (anonymous):

thats called completing the square